OpenStudy (anonymous):

Complete the following table, where information is given about a right triangle with legs a and b, hy-potenuse c, perimeter P, and area A. Round to the nearest hundredth...... a2.30 b_______ c_______ P 21.02 A 10.60

OpenStudy (anonymous):

How do I set it up?

sam (.sam.):

right triangle

sam (.sam.):

P=a+b+c 21.02=2.30+b+c...............(1) A=(1/2)(a)(b) 10.60=(1/2)(a)(b)..................(2) Find b when a= 2.30 then use (b) to (1) and solve for (c)